OpenStudy (anonymous):

Calculus help please: A tin can (right circular cylinder) with top and bottom is to have volume V . What dimensions (the radius of the bottom and the height) give the minimum total surface area?

OpenStudy (anonymous):

I am given both r and h, and am able to show that \[r=\sqrt[3]{V/(2 \pi)}\] , which is what is given

OpenStudy (anonymous):

ick make the diameter the height i guess we have to do a bunch of work, but that is the punch line

OpenStudy (anonymous):

V is fixed, it is a constant, treat it like one

OpenStudy (anonymous):

what you need is a formula for the surface area

OpenStudy (anonymous):

top and bottom have area \(\pi r^2\) for a total of \(2\pi r^2\)

OpenStudy (anonymous):

h is given as \[h=2\sqrt[3]{V/(2 \pi)}\] , how do I find that with the r I have found? I keep coming up with other answers when I substitute r in either the volume or area formula for the tin can

OpenStudy (anonymous):

area of the cylinder is \(2\pi rh\)

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, plus top and bottom

OpenStudy (anonymous):

total surface area is \[2\pi r^2+2\pi h\] the latter from the circumference of the circle times the height

OpenStudy (anonymous):

that is what you are trying to minimize

OpenStudy (anonymous):

Volume is fixed at \[V=\pi r^2h\]

OpenStudy (anonymous):

solve that not for \(r\) but for \(h\) and get \[h=\frac{V}{\pi r^2}\]

OpenStudy (anonymous):

yes, I'm told to minimize the surface area \[2\pi r^2+2\pi rh\]

OpenStudy (anonymous):

yes replace \(h\) by \(\frac{V}{\pi r^2}\)

OpenStudy (anonymous):

did that, took the derivative of the equation, using V as constant, set it to zero, and came up with the r= equation above

OpenStudy (anonymous):

then you get a function of \(r\) namely \[V(r)=2\pi r^2+\frac{V}{\pi r^2}\]

OpenStudy (anonymous):

the derivative is \[4\pi r-\frac{2V}{\pi r^3}\]

OpenStudy (anonymous):

ooh i think i made a mistake sorry

OpenStudy (anonymous):

so I have r (above) wondering why I keep getting a different h value than what the professor has given us as the correct answer (above)

OpenStudy (anonymous):

\[V=2\pi r^2+2\pi rh\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

put \(h=\frac{V}{2\pi r^2}\) and get \[V(r)=2\pi r^2+\frac{2V}{r}\]

OpenStudy (anonymous):

ok still screwing up, should be \(h=\frac{V}{\pi r^2}\) but \[V(r)=2\pi r^2+\frac{2V}{r}\] is right

OpenStudy (anonymous):

not the way I went with it

OpenStudy (anonymous):

taking derivatives gives \[V'(r)=4\pi r-\frac{2V}{r^2}\] yes?

OpenStudy (anonymous):

I took the derivative of the area fomula after solving the volume formula for h, and substituting that into the area formula

OpenStudy (anonymous):

you are trying to minimize the surface area, not the volume

OpenStudy (anonymous):

the volume is fixed, it is a number

OpenStudy (anonymous):

@xapproachesinfinity was helping with this same problem last night, I got to the point where I found r, but was unable to show steps to find the given h (above)

OpenStudy (anonymous):

oh i see you want r and h right?

OpenStudy (anonymous):

yes, vol is fixed, need to minimize surface area, both r and h are given above for optimized surface area with volume V, I need to show the steps. I can get to the point of showing what r is, but then when I go to substitute that into either the volume or surface area formula, I keep getting a different answer than what the professor has provided.

OpenStudy (xapproachesinfinity):

seems you r is correct just did it the h you did what i said yesterday correct and you didn't get a good formula matching what yu need

OpenStudy (anonymous):

\[4\pi r^3-2V=0\\ 4\pi r^3=2V\\ r^3=\frac{V}{2\pi}\\ r=\sqrt[3]{\frac{V}{2\pi}}\]

OpenStudy (anonymous):

h is given above, can you show me how to find THAT h, using the r as satellite has providded?

OpenStudy (anonymous):

correct satellite

OpenStudy (anonymous):

solve \[\pi r^2h=V\] for \(h\)

OpenStudy (anonymous):

h=V/(pir^2)

OpenStudy (anonymous):

using the r above

OpenStudy (anonymous):

that's where I get screwed up, something I'm doing isn't working out, can you show me your algebra?

OpenStudy (anonymous):

\[\pi(\sqrt[3]{\frac{V}{2\pi}})^2h=V\]\]

OpenStudy (xapproachesinfinity):

try this replace V by pi r^2h cancel and rewrite the radical to h=2r

OpenStudy (anonymous):

jeez this is hard to type

OpenStudy (anonymous):

that is a good idea again the punch line is the height is the diameter

OpenStudy (anonymous):

why is it that the height is the diameter?

OpenStudy (anonymous):

\[\pi(\sqrt[3]{\frac{V}{2\pi}})^2h=V\] we can use rational exponents i guess, maybe that will help

OpenStudy (xapproachesinfinity):

\[r=\sqrt[3]{\frac{\pi r^2h}{2\pi }}=\sqrt[3]{\frac{r^2h}{2}}\]

OpenStudy (xapproachesinfinity):

cube all sides to get \[r^3=\frac{r^2h}{2} \Longrightarrow h=2r\]

OpenStudy (anonymous):

X, how did you know to look for that, experience, or looking to get rid of V?

OpenStudy (anonymous):

\[h=V\times \left(\sqrt[3]{\frac{2\pi}{V}}\right)^2\frac{1}{\pi}\]

OpenStudy (anonymous):

grind it out if you like, i like @xapproachesinfinity way more

OpenStudy (anonymous):

yeah, tried grinding it out, and got a different answer than the prof...Satellite, any chance you'd grind it out so I can learn some algebra?

OpenStudy (xapproachesinfinity):

just after eyeballing, well i won't say i have enough experience :) if you tried replacing r that is good too

OpenStudy (anonymous):

guys, off to bed, thanks for all of the help, will check responses in the morning

OpenStudy (anonymous):

\[h=V\times V^{-\frac{2}{3}}\] etc etc